OpenStudy (anonymous):

Suppose that sin(theta)=3/5 and theta is in the second quadrant. Find sin(2(theta)), cos(2(theta)), and tan(2(theta)) exactly.

OpenStudy (anonymous):

use arcsin(3/5) to find theta

OpenStudy (anonymous):

i don't even know what arcsin(3/5) my instructor just does the problems while talking to himself, and i don't get it how he does them. That's why i'm always stuck for 10 hours only on my math homework.

OpenStudy (anonymous):

You have to use a calculator.

OpenStudy (anonymous):

Unless rouault knows of a unit circle measure that we can use. I'm going to have to wimp out on this one. Too tired. Sorry :/

OpenStudy (anonymous):

so sin2theta=2sin(theta)cos(theta)

OpenStudy (anonymous):

i assume this is Double-angle identities right?

OpenStudy (anonymous):

sin(theta)=3/5 theta=arcsin(3/5) sin(2arcsin(3/5)) double angle identity sin(2arcsin(3/5)) = 2sin(arcsin(3/5))(cos(arcsin(3/5)) and so on

OpenStudy (anonymous):

arcsin? isn't that the same thing as sin2x=2sin x cos x cos 2x=cos^2 x -sin^2 x =1-2 sin^2 x =2 cos^2 x-1 tan2x=(2tanx)/(1- tan^2 x)

OpenStudy (anonymous):

its inverse, instead of from angle to sides its from sides to angles

OpenStudy (anonymous):

arcsin means to ask the calculator what the degree was for sin. For example, (I'm making up numbers here), sin50 = 13/12 but arcsin 13/12 = 50 Note that those numbers are not factual.

OpenStudy (anonymous):

to solve this, instructor indicated to use Double-angle identities

OpenStudy (anonymous):

this is what i got \[\sin2\theta=\frac{-24}{5}\] \[\cos2\theta=\frac{7}{25}\] \[\tan2\theta=\frac{-24}{7}\]

OpenStudy (anonymous):

to solve cos(arcsin(3/5)) find out what the triangle is for arcsin(3/5) looks like. It is a 3 4 5 righttriangle in standard position. Find the cos of that tringle and you have the answer, 4/5

OpenStudy (anonymous):

sin(arcsin(theta)) will cancel since they are inverse of eachother sin(arcsin(theta))= theta

OpenStudy (anonymous):

the answer to the first one is 24/25

OpenStudy (anonymous):

or rather, using theta is sort of confusing, sin(arcsin(3/5))= 3/5

OpenStudy (anonymous):

Rouault thanks. I've got to move on. have a wonderful day.